OpenStudy (anonymous):

Sharyn worked as a Registered nurse in london last summer. The fisrt week she was there she worked a total of 23 hours at a city hospital. She was paid $13.50 and hour for a day shift and $16.50 an hour for night shift work. She earned $352.50 for her first week of work. How many hours did she spend on the day shift? how many hours did she spend on the night shift?

OpenStudy (anonymous):

13.50x+16.50y=352.5 x+y=23 y=23-x 13.50x+16.50(23-x)=352.5